OpenStudy (anonymous):

present value = 500 000 A bank offer interest rate of 9% p.a compounded quarterly for first 5 years. Rate after first 5 years is the reduce to 7.5% per year compounded monthly. Calculate new monthly payment after 5 years assuming company wants to keep total term of loan at 20 years.

OpenStudy (anonymous):

What level of math is this?

OpenStudy (anonymous):

Is no monthly payment made for the first 5 years?

OpenStudy (anonymous):

its financial math

OpenStudy (anonymous):

not sure im new to this

OpenStudy (anonymous):

There are infinite solutions if you allow two different monthly payments for each period.

OpenStudy (anonymous):

The remaining principal of a loan after \(t\) months is given by the equation: \[ P_t= P_0r^t-m\left(\frac{1-r^t}{1-r}\right) \]Where: \(P_0\) is the initial principal \(r\) is the monthly interest rate \(m\) is the monthly payment

OpenStudy (anonymous):

This equation can help you if you split it up.

OpenStudy (anonymous):

It can only be solved if we know the monthly payment for the first 5 years or know that it is the same. But they said 'new monthly payment' so I assume it is different.

OpenStudy (anonymous):

Maybe I'm misinterpreting this question. Doesn't anyone else understand the question better? Is there anyone out there?

OpenStudy (amistre64):

i read it as more of an adjustable rate loan determine the payments for a 25 year fixed rate, balance that out for 5 years to determine the remaining payment amount for the remianing 20 years

OpenStudy (amistre64):

if we use the wio formula syntax :) \[P_0r^{4t}~\frac{1-r}{1-r^{4t}}=m\] where r=1+.09/4

OpenStudy (amistre64):

run that for 4*5 periods to determine the balance to refinance ...

OpenStudy (anonymous):

i dont understand :(